OpenStudy (anonymous):

Which table of values can be used to graph the function. y = -4x + 3? A.) X Y -1 -1 0 3 1 7 2 11 B.) X Y -3 -9 -1 -1 1 7 3 15 C. X Y 0 3 1 -1 2 -5 3 -9 D.) X Y 0 3 1 7 2 11 3 15

jimthompson5910 (jim_thompson5910):

hint: plug in x = 1

jimthompson5910 (jim_thompson5910):

tell me what you get when you do

OpenStudy (anonymous):

x=1?

OpenStudy (shamil98):

y = -4x + 3 for when x = 1 y = -4(1) +3 evaluate it

jimthompson5910 (jim_thompson5910):

y = -4x + 3 y = -4(1) + 3 ... plug in x = 1 (ie replace x with 1) y = ???

OpenStudy (anonymous):

-4(1)**

OpenStudy (anonymous):

-1? correct?

jimthompson5910 (jim_thompson5910):

so if x = 1, then y is y = -1

jimthompson5910 (jim_thompson5910):

therefore, one row should be 1 -1

OpenStudy (anonymous):

y=-4(1)+3 y= -4 + 3

OpenStudy (shamil98):

now just look at which table has that and you have your answer

OpenStudy (anonymous):

there is no table for x = 1 and y= -1.....

jimthompson5910 (jim_thompson5910):

yes there is

OpenStudy (shamil98):

X Y 0 3 1 -1 2 -5 3 -9

jimthompson5910 (jim_thompson5910):

look for a table where it has 1 in the x column and -1 in the y column (for the same row)

OpenStudy (anonymous):

ohh ok sry:-/

OpenStudy (anonymous):

Thank you:-)